A supplier provides parts to a manufacturing company that demands JIT deliveries. At the present time it takes 6 hours to make a round-trip between the supplier's warehouse and the customer, including loading, travel and unloading time. The lot size is 12 pallet loads on a truck, and the manufacturer uses 2 pallets per hour. How many trucks are needed to ship the pallets to the manufacturer? What is likely to happen if the trucks break down? How can the supplier ensure that the customer does not run out of parts even in the face of delivery problems or other uncertainties? What will happen if the manufacturer runs into trouble and shuts down for a period of 6 hours? If unexpected overtime is required, how do you effectively communicate the need to your workforce?
Answer #1

a. N= DT/C = 2(6)/12=1truck

2. On the off chance that the truck separates the provider won't have the capacity to convey the important beds n time. Hence, one extra truck is most likely expected to safeguard conveyance.

3. Aside from having one more truck to convey in the event of crisis, the provider may wish to have one day request security stock at the client's stockroom, which is equivalent to (2 beds for each hour) x (working hours) This will ensure against different possibilities, for example, provider plant shutdown or deficiencies for a brief span, shutdown of the interstate for climate or mischances, and so forth

4. In the event that the maker closes down for six hours, the providers' conveyance cycle will be aggravated.

5.communication should be done via any relevant techniques put in place in order to ensure continuity in the production process